A coffee merchant combines coffee that costs $7 per pound with coffee that costs $2.60 per pound. How many pounds of each should be used to make 22 lb of a blend costing $5.90 per pound?
Answer by lwsshak3(11628)

A coffee merchant combines coffee that costs $7 per pound with coffee that costs $2.60 per pound. How many pounds of each should be used to make 22 lb of a blend costing $5.90 per pound?
***
let x=amt of $7 coffee used
22-x=amt of $2.60 coffee used
..
7x+2.60(22-x)=5.90*22
7x+57.20-2.60x=129.80
4.40x=72.60
x=16.5
22-x=5.5
amt of $7 coffee used=16.5 lbs
amt of $2.60 coffee used=5.5 lbs